OpenStudy (anonymous):

Select the true statements about half-angle identities. a)The half-angle identity for cosine is cosA/2 = square root of (1 - cos2A)/2. b)The half-angle identity for sine is sinA/2 = 2sinAcosA. c)The half-angle identity for tangent is tanA/2 = 2cos2A - 1. d)Half-angle identities of cosine, sine, and tangent are in terms of cosine of the whole angle.

OpenStudy (anonymous):

b is false since \(\sin(2A)=2\sin(A)\cos(A)\)

OpenStudy (anonymous):

okay thank u :)

OpenStudy (anonymous):

but do u know which is true?

OpenStudy (anonymous):

here is a nice summary of the half angle formulas

OpenStudy (anonymous):

http://www.freemathhelp.com/trig-double-angles.html

OpenStudy (anonymous):

you can see that the first one is wrong as well

OpenStudy (anonymous):

so the answer is d? and are u talking about a being wrong

OpenStudy (anonymous):

actually i do not see one what is right hmmm

OpenStudy (anonymous):

huh?? lol

OpenStudy (anonymous):

the half angle formulas for tangent are in terms of sine and cosine, as you can see

OpenStudy (anonymous):

so the last one is wrong

OpenStudy (anonymous):

oh no scratch that, the last one is right

OpenStudy (anonymous):

ohhh okay soo d :) lol thsnks

OpenStudy (anonymous):

you can write \[\tan(\frac{A}{2})=\pm\sqrt{\frac{1-\cos(A)}{1+\cos(A)}}\]

OpenStudy (anonymous):

yes it is d